# Mapping and Scheduling Spiking Neural Networks On Segmented Ladder Bus Architectures

### v1

@@ 2 $RoutedList=\emptyset$;
4       $Path=A\_star(i,j)$; /* find the path for the link using A* */
@@ 4.1 Benchmarks
fashion-mnist [25] 24 5.33 0.46 5
@@
emnist [26] 30 5.37 0.37 6
ResNet [27] 96 11.13 0.23 10
@@ 5.3 Comparison to Other Architectures
In this section, we compare the latency, energy consumption, and energy-delay product (EDP) of four architectures: mesh-based NoC, butterfly NoC, NeuSB, and MASS-SR.
@@
Figure 12 shows that MASS-SR has a slightly higher latency compared to other architectures due to spike traffic scheduling. On average, MASS-SR experiences a 1.24 times increase in latency compared to Mesh NoC, with a maximum increase of 1.7 times. However, this latency is still negligible for the streaming application domain which we target, where the sample periods are above 1 msec. NeuSB achieves the lowest latency due to its dedicated bus lane per master, which reduces routing complexity and delays.

### v2

@@ 2 $RoutedList=\emptyset$;
4  Path=A_star(i,j)Path=A\_star(i,j); /* find the path for the link using A* */
@@ 4.1 Benchmarks
The hardware implementation details of the segmented ladder bus were presented in our prior work [huynh2024adiona]. In that work, we implemented both the 3-way segmented switches and a standard NoC router on an FPGA to evaluate their latency and power characteristics. These measured values were then used to calibrate our cycle-accurate simulator extended from Noxim [catania2015noxim], enabling it to model the segmented ladder bus alongside NoC variations with realistic energy and performance metrics. All evaluations are based on this extended simulator, which incorporates FPGA-measured dynamic energy and latency per hop and per switch configuration.
@@
fashion-mnist [xiao2017fashion] 24 5.33 0.46 5
@@
emnist [cohen2017emnist] 30 5.37 0.37 6
ResNet [he2016deep] 96 11.13 0.23 10
@@ 5.1 Cluster mapping
Reducing spike loss is critical in spiking neural networks, as each spike encodes meaningful temporal information that contributes to downstream neuronal activations. Lost spikes, particularly in later or sparsely connected layers, can lead to insufficient membrane potential buildup, delayed firing, or complete suppression of neuron activity, which may ultimately result in degraded inference accuracy. By reducing spike loss during communication, our framework ensures that application-level functionality and accuracy are preserved.
@@ 5.3 Comparison to Other Architectures
In this section, we compare the latency, energy consumption, and energy-delay product (EDP) of four architectures: mesh-based NoC [truenorth], butterfly NoC [jiang2021butterfly], NeuSB [balaji2022NeuSB], and MASS-SR.
@@
Figure 12 shows that MASS-SR has a slightly higher latency compared to other architectures due to spike traffic scheduling. On average, MASS-SR experiences a 1.24 times increase in latency compared to Mesh NoC, with a maximum increase of 1.7 times. When the application has a high number of spikes per connection (e.g., the LeNet benchmark), fewer scheduling delays are required, resulting in lower latency. In contrast, for applications with more uniformly distributed spike traffic, the scheduler must serialize more connections to avoid path conflicts, which naturally leads to increased latency. However, this latency is still negligible for the streaming application domain which we target, where the sample periods are above 1 msec. NeuSB achieves the lowest latency due to its dedicated bus lane per master, which reduces routing complexity and delays.
@@ Acknowledgment
Phu Khanh Huynh received a Ph.D. degree from the Department of Electrical and Computer Engineering, Drexel University, Philadelphia, PA in 2025. His research interests include designing, exploration, and optimization of neuromorphic computing systems, with focus on spiking neural networks (SNN).
Francky Catthoor received the engineering degree and a Ph.D. in electrical engineering from the Katholieke Universiteit Leuven, Belgium, in 1982 and 1987 respectively. Between 1987 and 2000, he headed several research domains in the area of high-level and system synthesis techniques and architectural methodologies at IMEC. Since 2000 he has also been strongly involved in other activities including co-exploration of application, computer architecture and deep submicron technology aspects, biomedical systems and IoT sensor nodes, and photovoltaic modules combined with renewable energy systems, all at the Inter-university Micro-Electronics Center (IMEC), Heverlee, Belgium. After his official retirement from IMEC in Oct 2024, he is now affiliated with the Microlab of NTUAthens in Greece as an associated guest professor. He is emeritus full professor at the EE department of the K.U.Leuven. He was elected an IEEE Fellow in 2005.
Anup Das is an Associate Professor at Drexel University and Associate Department Head for Graduate Studies. He received a Ph.D. in Embedded Systems from the National University of Singapore in 2014. Following his Ph.D., he was a postdoctoral fellow at the University of Southampton and a researcher at IMEC. His research focuses on neuromorphic computing and architectural exploration. He received the United States National Science Foundation CAREER Award in 2020 and the Department of Energy CAREER Award in 2021 to investigate the reliability and security of neuromorphic hardware. He is a senior member of the IEEE and a member of the ACM.